Fermi National Laboratory
 

Operation Summary for July 9, 2010 - July 12, 2010

 

The Integrated Luminosity for the period from 7/5/10 to 7/12/10 was 51.16 pb-1.
NuMI reported receiving 4.89E18 protons on target during this same period.

Friday July 9

The day shift began with Operations monitoring the Tevatron (TeV) store 7960 (92.86E30) and the Recycler stash (220.1E10), with the Antiproton Source (Pbar) down and not stacking (1.9E10) due to a bend bus power supply problem, with the Linac, Booster and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, power supply technicians had the Pbar Debuncher bend bus power supply repaired and back online by 11:00 AM; stacking resumed. Operations terminated store 7960 at 1:15 PM. TeV personnel began conducting studies that lasted the rest of the shift.

During the evening shift, Operations established HEP store 7964 at 6 PM with an average initial luminosity (AIL) of 297E30. A MI Lambertson power supply (LAM10) tripped off due to a transformer over-temperature indication. EE Support personnel came in and found that two of the four fans in the supply had failed.

Saturday July 10

The midnight shift began with Operations monitoring store 7964 (120.24E30) and the stash (99.5E10), with Pbar stacking (20.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-977 at 4:15 AM. Operators accessed MI-31 to investigate four corrector magnet power supplies that had tripped off. They checked the fuses on the power supply chassis that control the correctors; reseating the fuses corrected the chassis issue.

During the day shift, Operations terminated store 7964 at 1:15 PM and put the TeV into shot setup and established store 7965 at 3:15 PM with and AIL of 326E30. At 7:45 PM, store 7965 aborted when the ramp dumped without quenching. Operators investigated the cause.

During the evening shift, a TeV person came in and conferred with EE Support about the abort. They determined that the TeV sector C2 safety Coordinator was the cause of the abort. An EE Support technician came in and replaced the Safety Coordinator. Operations had the TeV up and ramping by 11:50 PM.

Sunday July 11

The midnight shift began with Operations ramping the TeV, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into shot setup at 2:20 AM and established store 7969 at 3:20 AM with an AIL of 259E30. A number of Electron Cooling (ECool) corrector magnet power supplies (R:CYC30I, R:CXC30I, R:CYC39I, and R:CXC39I) tripped off just as they did on the previous midnight shift. Operators called in an EE Support technician to investigate and the tech changed out the faulty chassis. T-977 started taking beam at 4:49 AM.

During the day shift, the 400 MeV area tripped off a number of times due to tune problems; operators informed Booster personnel.

During the evening shift, Operations terminated store 7969 at 6:20 PM. Operations terminated the store early due to a necessary Pelletron access. Operations put the TeV into shot setup and established store 7970 at 7:45 PM with an AIL of 170E30. Booster personnel investigated several issues with the 400 MeV chopper system; they made repairs and Operations had beam re-established by 6 PM. Operators reported that the Pelletron developed voltage regulation problems; ECool personnel investigated and traced the problem back to a triode inside the tank. Operations called in Mechanical Support personnel to prepare the tank for access.

Monday July 12

The midnight shift began with Operations monitoring store 7970 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the NuMI horn power supply tripped off at 3:40 AM, which was 20 minutes prior to the scheduled shutdown. NuMI personnel reported a large LCW leak coming from the horn power supply. Operators isolated the leak. Operations established beam to T-977 at 4:30 AM. Operations terminated store 7970 at 6:05 AM, put the TeV into shot setup, and established store 7971 at 7:25 AM with an AIL of 131E30.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow the following work: Preaccelerator personnel will switch from the I- Source to the H- Source. Recycler personnel will conduct studies and work on the Pelletron. MI personnel will conduct studies. On Tuesday at 6 AM, MTA will take 10 pulses of beam and then go into a six-hour access that will halt Linac beam.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/12/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory